### v3

# Constraints on the Variation of the QCD Interaction Scale $\Lambda_{\text{QCD}}$

Abstract
Laboratory and astrophysical tests of “constant variation” have so far concentrated on the dimensionless fine-structure constant $\alpha$ and on the electron or quark mass ratios $X_{e,q}=m_{e,q}/\Lambda_{\text{QCD}}$, treating the QCD scale $\Lambda_{\text{QCD}}$ as unchangeable. Certain beyond Standard Model frameworks, most notably those with a dark matter or dark energy scalar field $\phi$ coupling with the gluon field, would make $\Lambda_{\text{QCD}}$ itself time dependent while leaving $\alpha$ and the electron mass untouched. Under the minimal assumption that this gluonic channel is the sole $\phi$ interaction, we recast state-of-the-art atomic clock comparisons into $\dot{\Lambda}_{\text{QCD}}/\Lambda_{\text{QCD}}=(3.2\pm 3.5)\times 10^{-17}\ \text{yr}^{-1}$ limits, translate the isotope yields of the 1.8-Gyr-old Oklo natural reactor into a complementary geophysical limit of $|\delta\Lambda_{\text{QCD}}/\Lambda_{\text{QCD}}|<2\times 10^{-9}$ over that time span, corresponding to the linear drift limit $|\dot{\Lambda}_{\text{QCD}}/\Lambda_{\text{QCD}}|<1\times 10^{-18}\text{yr}^{-1}$, and show that the proposed $8.4$ eV 229Th nuclear clock would amplify a putative $\Lambda_{\text{QCD}}$ drift by four orders of magnitude compared with present atomic clocks. We also obtain constraints from quasar absorption spectra and Big Bang Nucleosynthesis data.
Introduction
The search for the variation of fundamental constants is a well-established area of research motivated by theories beyond the Standard model that unify gravity with the other fundamental forces, as well as cosmological models involving scalar fields which predict a space-time variation of these constants. Constraints on such variation have been derived from Big Bang Nucleosynthesis (BBN), quasar absorption spectra, the cosmic microwave background (CMB) and the Oklo natural nuclear reactor, see e.g. the review [1] and references therein. The stringent bounds on the linear drift of the fundamental constants have been obtained using measurements of the temporal drift of atomic clock transition frequencies. These experiments measure the variation of the dimensionless ratio of two transition frequencies $\nu_{1}$ and $\nu_{2}$
$\delta\left(\frac{\nu_{1}}{\nu_{2}}\right)\,$, (1)
which have different dependencies on the fundamental constants. As a result, these measurements may be used to place constraints on the variation of these constants. Experimental results have been obtained by measuring the oscillating frequency ratios of electron transitions in a range of systems, including Dy/Cs [2], Rb/Cs [3], Yb/Cs [4], Sr/H/Si cavity [5], Cs/cavity [6], Cs/H [7], Al+/Hg+ [8] and Yb+/Yb+/Sr [9, 10]. The results of these experiments have been used to place constraints on the variations of fundamental constants using the sensitivity calculations performed in Refs. [11, 12, 13, 14, 15, 16, 17] and several other papers referenced there.
It is important to note that if the absolute measurement of a transition frequency is performed, the dimensionless ratio of this frequency to the Cs hyperfine transition frequency, which defines the units of frequency and time, is being measured. The Cs transition frequency has a complicated dependence on the fundamental constants [15]. This principle applies to any measurement of variations of fundamental constants in any field of physics and astrophysics. Any measurement of a dimensionful parameter is, in fact, a measurement of the dimensionless ratio of that parameter to the unit used to measure it, which itself depends on the fundamental constants.
Because any measurable quantity is ultimately a dimensionless ratio, it can always be written solely in terms of dimensionless fundamental constants. To do so, one first identifies the particular dimensionless constants that enter a given experiment. By contrast, many papers discuss variations of dimensionful parameters, such as the speed of light $c$, Planck’s constant $\hbar$, Newton’s gravitational constant $G_{N}$, the QCD scale $\Lambda_{\text{QCD}}$, or particle masses. The problem is that the numerical values of these quantities depend on the system of units: in natural units, for example, $c=1$ by definition, so it cannot ”vary.” Any apparent change can be absorbed by a re-definition of units. Consequently, interpreting experimental results in terms of changing dimensionful constants is not model-independent; one must instead identify a set of dimensionless fundamental constants relevant to this measurement and calculate the dependence of an observed effect on these constants  [^1] It was noted in Refs. [18, 19], that the number of possible ”dimensionless” constants may increase when we allow for fast variations of the constants, where ”fast” is determined by the time scale of the response of the studied species or experimental apparatus used. In this case, the relevant dimensionless quantity is, for example, the ratio $m_{e}/\expectationvalue{m_{e}}$, and $\expectationvalue{m_{e}}$ is the time average. In this sense, one may say that the experimental signal depends on the variation of dimensionful constants (the electron mass $m_{e}$ in this example). Sensitivity to such variations requires two systems, one which has a faster response (such as an atom) and another which is more inertial (such as a cavity).
That said, fundamental Lagrangians are often written with dimensionful constants. To be consistent, one must include every relevant term in the Lagrangian when deriving how the dimensionless ratios of fundamental constants vary for a given measurement.
Alternatively, one may speak about the apparent variation of the fundamental constants, which may be just a convenient way to calculate the effects of specific terms in the interaction of Standard model particles with a dark matter (or dark energy [20, 21, 22, 23, 1]) scalar or pseudoscalar field $\phi$. For example, the interaction of fermions with $\phi$, which may be presented in the form $g\phi^{n}m\bar{f}f$, is similar to the fermion mass term $m\bar{f}f$. This means that one may define a variable fermion effective mass $m_{eff}=m(1+g\phi^{n})$, which depends on the field $\phi$. Here, $n=1,2$ for scalars and $n=2$ for pseudoscalars. Similarly, the interaction of $\phi$ with the electromagnetic field may be incorporated as a variable fine structure constant $\alpha_{eff}=\alpha(1+g\phi^{n})$. The effects of the variation of fermion masses and the variation of $\alpha$ have been calculated; see, e.g. [11, 12, 13, 24, 25, 14, 15, 16, 17]. Note that in this approach, the variation of measurement units is still a relevant issue if one considers measurements of dimensionful parameters. We note that using experimental results to place constraints on the variation of dimensionful parameters is only possible within a specific model of variation, e.g. an interaction with an ultralight dark matter field or dark energy field, which varies on a cosmological scale.
For clarity, we note that probing an extremely small fractional change in a constant does not require knowing that constant’s absolute value with comparable precision. What matters is the sensitivity coefficient that links observable shifts to fractional changes. For example, if an atomic transition frequency $\nu$ obeys $\delta\nu/\nu=k_{\alpha}(\delta\alpha/\alpha)$ one only needs a calculation of $k_{\alpha}$ and high precision measurement of $\delta\nu/\nu$. At present $\delta\alpha/\alpha$ is constrained to better than $10^{-18}$ per year, even though the absolute value of $\alpha$ is known only to $1.5\times 10^{-10}$.
The model-independent variations of dimensionless constants can be attributed to the fractional variation of three fundamental independent dimensionless parameters: the fine structure constant $\alpha$ and the ratios of the electron/quark masses to the QCD scale $\Lambda_{\text{QCD}}$, $X_{e}$ and $X_{q}$, where $\Lambda_{\text{QCD}}$ is defined as the position of the Landau pole in the logarithm for the running strong coupling constant, $\alpha_{s}(r)\sim\text{const}/\ln(\Lambda_{\text{QCD}}r/\hbar c)$.
We seek to examine the sensitivity of various systems to variations of the QCD scale $\Lambda_{\text{QCD}}$. Such a sensitivity appears in the frequency comparison of two microwave atomic clocks or optical and microwave clocks (see below). It is usually assumed that the comparison of two optical clock transitions is sensitive to variations of the fine structure constant only; however, following Ref. [9], we argue that this system is also sensitive to variations of $\Lambda_{\text{QCD}}$ and repurpose existing results to place constraints. We also examine the sensitivity of the fractional variation of the transition frequency between the ground state and the low-energy isomer in 229Th to variations in the QCD scale $\Lambda_{\text{QCD}}$, which appears to be enhanced relative to the proposed nuclear clock’s sensitivity to the quark mass $m_{q}$. Additionally, we place constraints on $\Lambda_{\text{QCD}}$ from quasar absorption spectra and Big Bang Nucleosynthesis. Finally, we reinterpret existing constraints from the Oklo natural nuclear reactor to place constraints on the variation of $\Lambda_{\text{QCD}}$.
I Constraints on $\Lambda_{\text{QCD}}$ via the Reinterpretation of Clock constraints
The experimental observation of variation in nature may be interpreted in a model-independent manner as the variation of at least one of three independent fundamental dimensionless ratios;
$\frac{\delta\alpha}{\alpha}\,,\ \frac{\delta X_{q}}{X_{q}}\,,\ \text{and}\ \frac{\delta X_{e}}{X_{e}}\,$, (2)
where $X_{q,e}$ corresponds to the ratio of the quark/electron mass to the QCD scale $\Lambda_{\text{QCD}}$. Here, we are working in natural units, setting $c=1$ and assuming that $\Lambda_{\text{QCD}}$ has dimension of mass. For example, the variation of the ratio of two atomic, molecular or nuclear transition frequencies may be presented as
$\frac{\delta(\nu_{1}/\nu_{2})}{\nu_{1}/\nu_{2}}=k_{\alpha}\frac{\delta\alpha}{\alpha}+k_{q}\frac{\delta X_{q}}{X_{q}}+k_{e}\frac{\delta X_{e}}{X_{e}}\,$, (3)
Table 1 lists the current constraints on the variation of the ratio of the two hyperfine transition frequencies in 87Rb and 133Cs and the constraints on the variation in the ratio of optical transition frequencies in 1H, 199Hg+, 171Yb+ and 87Sr to the hyperfine transition in 133Cs. In the previous calculations in Refs. [14, 15], the dependence on $X_{q}$ was calculated assuming that $\Lambda_{\text{QCD}}$ is constant and only the quark masses vary. Below, we will demonstrate that there is no need for such an assumption.
The dependence of the atomic hyperfine transition frequency on $X_{q}$ appears mainly from the dependence on the nuclear magnetic moment. The dependence of the nuclear magnetic moments on $m_{q}$ was calculated in Refs. [14, 15]. A less significant contribution to the $m_{q}$ dependence, which originates from the variation of the nuclear radius, was calculated in Refs. [26, 27].
The nuclear magnetic moment and nuclear radius are also sensitive to the QCD scale $\Lambda_{\text{QCD}}$. In effective field theory, the nucleon magnetic moments receive loop-level corrections from pion-nucleon interactions governed by the spontaneously broken chiral symmetry of QCD. These corrections provide a natural mechanism by which these magnetic moments can acquire a dependence on both the quark mass and $\Lambda_{\text{QCD}}$. The nucleon magnetic moment including such corrections can be schematically written as [28, 29]:
$\mu_{N}=\mu_{N}^{(0)}+\mu_{N,\pi}^{(1)}\,$, (4)
where $\mu_{N}^{(0)}$ denotes the static low-energy constant contributions at tree-level and $\mu_{N,\pi}^{(1)}$ represents the dominant corrections to the nucleon magnetic moments, measured in units of the nuclear magneton $e\hbar/2m_{p}c$, which at leading one-loop order are proportional to the pion mass $m_{\pi}$ [30, 31]:
$\mu_{N,\pi}^{(1)}\sim\frac{g_{A}^{2}m_{\pi}m_{N}}{8\pi f_{\pi}^{2}}\,$, (5)
where $g_{A}$ is the axial coupling, $m_{N}$ the nucleon mass, and $f_{\pi}$ the pion decay constant. The sign of this correction is positive for the proton magnetic moment and negative for the neutron magnetic moment. Since $m_{\pi}^{2}\sim m_{q}\Lambda_{\text{QCD}}$ via the Gell-Mann–Oakes–Renner relation  [32] and $f_{\pi}\propto\Lambda_{\text{QCD}}$, the pion mass inherits a dependence on the light quark mass $m_{q}$ and $\Lambda_{\text{QCD}}$. Consequently, the one-loop correction scales as:
$\mu_{N,\pi}^{(1)}\propto\left(\frac{m_{q}}{\Lambda_{\text{QCD}}}\right)^{1/2}\,.$ (6)
Thus, we see that nucleon magnetic moments are naturally sensitive probes of variations of dimensionless ratio of the quark mass $m_{q}$ and $\Lambda_{\text{QCD}}$.
Clock 1 Clock 2 Constraint (yr-1) Ref. Constants dependence
87Rb 133 Cs $(-0.5\pm 5.3)\times 10^{-16}$ [33] $\alpha^{-0.49}X_{q}^{-0.025}$
1H 133Cs $(-32\pm 63)\times 10^{-16}$ [34] $\alpha^{2.83}X_{q}^{-0.039}X_{e}$
199Hg+ 133Cs $(3.7\pm 3.9)\times 10^{-16}$ [35] $\alpha^{6.05}X_{q}^{-0.039}X_{e}$
171Yb+ 133Cs $(-3.1\pm 3.4)\times 10^{-17}$ [36] $\alpha^{1.93}X_{q}^{-0.039}X_{e}$
87Sr 133Cs $(-1.0\pm 1.8)\times 10^{-15}$ [37] $\alpha^{2.77}X_{q}^{-0.039}X_{e}$
Table 1: Constraints on the variation of the ratio of clock frequencies.
[…]
Under the minimal assumption that this gluonic channel is the sole $\phi$ interaction, the constraints presented in Table 1 can be translated into limits on the temporal variation of the QCD scale, $\Lambda_{\text{QCD}}$, via the dependence of each transition on $X_{q}$ and $X_{e}$. The dominant contribution to the sensitivity arises from the $X_{e}$ term, and the tightest constraints may be obtained from the 171Yb+/Cs clock measurement [36]:
$\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}=(3.2\pm 3.5)\times 10^{-17}\ \text{yr}^{-1}\,.$ (7)
I.1 Sensitivity of the comparison of two optical clock transition frequencies to $X_{e}$
Different types of atomic transition frequencies have a differing dependence on the fundamental constants. As discussed above, constraints obtained from the measurement of the ratio of optical and hyperfine transition frequencies have sensitivity to $X_{e}$ via a dependence on the electron to proton mass ratio $m_{e}/m_{p}$ and sensitivity to $X_{q}$ via dependence on the nuclear magnetic g-factor.
Following Ref. [9], we aim to propose a similar dependence on $X_{e}$ for the ratio of two optical frequencies, by probing the contributions of nuclear mass $m_{N}$ and nuclear radius $r_{N}$ to atomic transition frequencies (these contributions produce mass and field isotope shifts). These contributions are small when compared to the optical transition frequency; however, this mechanism indicates a sensitivity to the parameter $X_{e}=m_{e}/\Lambda_{\text{QCD}}$ from measurements comparing the ratio of two optical transition frequencies. In general, comparing two transition frequencies $\nu_{a}$ and $\nu_{b}$, we obtain
$\frac{\delta(\nu_{a}/\nu_{b})}{(\nu_{a}/\nu_{b})}=(K_{1}-K_{2})\frac{\delta r_{N}^{2}}{r_{N}^{2}}+(K_{3}-K_{4})\frac{\delta m_{N}}{m_{N}}\,$, (8)
where
$K_{1}$ $=\frac{K_{\text{FS}}^{\nu_{a}}r_{N,a}^{2}}{\nu_{a}}\,$, (9)
$K_{2}$ $=\frac{K_{\text{FS}}^{\nu_{b}}r_{N,b}^{2}}{\nu_{b}}\,$, (10)
$K_{3}$ $=\frac{K_{\text{MS}}^{\nu_{a}}}{\nu_{a}m_{N,a}}\,$, (11)
$K_{4}$ $=\frac{K_{\text{MS}}^{\nu_{b}}}{\nu_{b}m_{N,b}}\,$, (12)
where $K_{\text{MS}}$ and $K_{\text{FS}}$ are the mass and field shift coefficients respectively and $m_{N}$ is the mass of an atom with atomic mass number $A$ (which is largely determined by the nuclear mass $m_{N}$). The mass shift term dominates for light nuclei whilst the field shift term dominates for heavy nuclei. Results have been obtained for the ratio of the optical transition frequencies of $^{1}S_{0}\rightarrow\ ^{3}P_{0}$ transitions in Al+ and $^{2}S_{1/2}\rightarrow\ ^{2}D_{5/2}$ transitions in Hg+ [38, 39, 40] and the ratio of the $^{2}S_{1/2}(F=0)\leftrightarrow\ ^{2}F_{7/2}(F=3)$ electric octupole (E3) and $^{2}S_{1/2}(F=0)\leftrightarrow\ ^{2}D_{3/2}(F=2)$ electric quadrupole (E2) transition frequencies in Yb+ [9]. We have previously used Equation (8) along with these results to place constraints on the variation of hadronic parameters, see Refs. [41, 40].
When considering low-mass nuclei, the sensitivity to the parameter $X_{e}$ may be yielded from the dependence of optical transition frequencies on the reduced mass (the normal mass shift)
$\bar{\mu}=\frac{m_{e}m_{N}}{m_{e}+m_{N}}\,$, (13)
where the nuclear mass $m_{N}$ is proportional to the nucleon mass, $m_{N}\approx Am_{n}$. The sensitivity to other hadronic parameters comes about from the energy shift due to a finite nuclear size.
These effects are extracted from the measurement of the variation of the ratio of two atomic transition frequencies which have different dependencies on varying parameters. In this case, one may say that the dimensionless parameters appear after dividing the variations of dimensionful parameters by the appropriate atomic units of energy $E_{h}=\alpha^{2}m_{e}c^{2}$, length $a_{B}=\hbar/(mc\alpha)$, etc.
In high-mass nuclei, the relative change in energy (or electronic transition frequency $\nu$) in atoms is proportional to
$\frac{\delta E}{E}\propto\left(\frac{r_{N}}{a_{B}}\right)^{2}=\left(\frac{r_{N}m_{e}\alpha c}{\hbar}\right)^{2}\,.$ (14)
Setting $c=\hbar=1$, and noting the fact that the nuclear radii in all nuclei are reasonably described by a universal formula $r_{N}=A^{1/3}r_{0}$, where $r_{0}\approx 1.2\ \text{fm}$ is the internucleon distance, we yield
$\left(\frac{r_{N}}{a_{B}}\right)^{2}=\left(r_{0}A^{1/3}m_{e}\alpha\right)^{2}\,.$ (15)
The distance between the nucleons $r_{0}$ is determined by the minimum of the potential in the nucleon-nucleon interaction. In the Walecka model [42] this potential is produced by a vector meson exchange repulsion at short distances and sigma meson exchange attraction at larger distances. Both the vector meson and sigma meson masses are proportional to $\Lambda_{\text{QCD}}$, so the range of these potentials is inversely proportional to $\Lambda_{\text{QCD}}$. Thus $r_{0}\propto\frac{1}{\Lambda_{\text{QCD}}}$ and
$\left(\frac{r_{N}}{a_{B}}\right)^{2}$ $\propto\left(\frac{A^{1/3}m_{e}\alpha}{\Lambda_{\text{QCD}}}\right)^{2}$ (16)
Assuming independent variations, the constraints on the variation of the nuclear charge radius $r_{N}$ presented in these references may be repurposed into constraints on the parameter $X_{e}$, noting that the variation of the $r_{N}^{2}$ is proportional to (see Equation (16))
$\frac{\delta(r_{N}/a_{B})^{2}}{(r_{N}/a_{B})^{2}}\sim 2\left(\frac{\delta m_{e}}{m_{e}}+\frac{\delta\alpha}{\alpha}-\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}\right)\sim 2\left(\frac{\delta X_{e}}{X_{e}}+\frac{\delta\alpha}{\alpha}\right)\,.$ (17)
The numerical constraints obtained from the systems of interest are presented in Table 2.
Clock 1 Clock 2 Constraint (yr-1) Ref. $\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}$ (yr-1)
Al+ Hg+ $(-5.3\pm 7.9)\times 10^{-17}$ [8] $(5.5\pm 8.5)\times 10^{-15}$
Yb+(E3) Yb+(E2) $(-1.2\pm 1.8)\times 10^{-18}$ [10] $(0.25\pm 0.38)\times 10^{-15}$
Table 2: Constraints on fractional variation of $\Lambda_{\text{QCD}}$ from the comparison of two optical clock transition frequencies in different systems.
II Sensitivity of the 229Th based nuclear clock to $\Lambda_{\text{QCD}}$
Refs. [43, 26] have calculated the sensitivity of the binding energy and nuclear radius in various light nuclei to the different hadron masses. These calculations were extrapolated to heavier nuclei. In such nuclei, the errors produced by such extrapolation were found to be smaller than the errors of direct calculations in heavy nuclei. Assuming that $\Lambda_{\text{QCD}}$ does not vary, these calculations were used to determine that the effects of the variation of the quark mass $m_{q}$ are enhanced roughly $10^{5}$ times in the $8.4$ eV “nuclear clock” transition between the ground and first excited states in 229Th and about $10^{8}$ times in the relative shift of the 0.1 eV compound resonance in 150Sm.
Let us now investigate the implications of interpreting their results as being due to a variation of $\Lambda_{\text{QCD}}$. Using Table III of Ref. [26], the variation of the binding energy $E_{S}$ is related to the variation of the hadron masses by
$\frac{\delta E_{S}}{E_{S}}\approx\Delta\mathcal{E}(m_{N})\frac{\delta m_{N}}{m_{N}}+\Delta\mathcal{E}(m_{\pi})\frac{\delta m_{\pi}}{m_{\pi}}\,$, (18)
where $m_{N,\pi}$ represent the nucleon and pion masses and $\Delta\mathcal{E}(m_{N,\pi})$ represent their respective dimensionless derivatives
$\Delta\mathcal{E}(m_{N,\pi})=\frac{\delta E_{S}/E_{S}}{\delta m_{N,\pi}/m_{N,\pi}}\,$, (19)
which were roughly calculated to be $\Delta\mathcal{E}(m_{N})\approx 12$ and $\Delta\mathcal{E}(m_{\pi})\approx-6$.
The variation of the nucleon mass $m_{N}$ can be related to the variation of the fundamental constants by [14, 44, 45, 46, 47, 48]
[…]
where $m_{q,-}=(m_{u}-m_{d})$ is the mass difference. The dependence of the variation of the nucleon mass on this parameter is weak, and we may neglect it. We will also assume that $\frac{\delta m_{s}}{m_{s}}=\frac{\delta m_{q}}{m_{q}}$, where $m_{s}$ is the mass of the strange quark (this assumption may be motivated by the Higgs mechanism of mass generation). Noting the dependence of the pion mass on $\Lambda_{\text{QCD}}$ and $m_{q}$ from the Gell-Mann-Oakes-Renner relation, $m_{\pi}^{2}\sim\Lambda_{\text{QCD}}m_{q}$ [32], we yield the following result for the variation of the binding energy $E_{S}$
$\frac{\delta E_{S}}{E_{S}}\approx 7.9\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}-1.5\frac{\delta m_{q}}{m_{q}}\,.$ (21)
This expression indicates that the variation of the binding energy $E_{S}$ seems to be $\sim 5$ times more sensitive to variations of $\Lambda_{\text{QCD}}$ than to variations of the average light quark mass $m_{q}$.
Note that the binding energy $E_{S}$ includes contributions from the strong interaction and the kinetic energy. The result for $\delta E_{S}/E_{S}$ was extrapolated from calculations performed for light nuclei, since the binding energy per nucleon is approximately the same in light and heavy nuclei. However, this expression does not include the contribution of the electromagnetic energy $E_{C}$ which is small in light nuclei ($E_{C}\propto Z(Z-1)$).
It may be instructive to estimate separately kinetic energy contribution $E_{K}\propto 1/m_{n}r_{N}^{2}$. Using Equation (20), together with $r_{N}=A^{1/3}r_{0}$ and the scaling $r_{0}\propto\Lambda_{\text{QCD}}^{-1}$, the fractional variation is
[…]
implying that in the kinetic energy the $\Lambda_{\text{QCD}}$ term strongly dominates over the $m_{q}$ term.
These results have implications for experiments with the “nuclear clock” transition between the ground and first excited states of the 229Th nucleus, first proposed in Ref. [49], where effects of the variation of fundamental constants are strongly enhanced [50]. This transition has been measured by a range of experiments, see e.g. [51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. The most precise measurement gives a transition frequency of $\omega=2,020,407,384,335\ (2)$ kHz ( 8.36 eV) for 229Th dopant ions embedded in a calcium fluoride crystal [60].
Using the results presented above, we may write this transition frequency $\omega$ as a sum of nearly canceling components that have different dependencies on the fundamental constants. The low-lying isomeric state in 229Th is theorised to exhibit a different shape and spin configuration to the ground state [61, 62], resulting in a small energy shift due to the residual strong and electromagnetic effects. As such, we may write $\omega$ as a sum of the electromagnetic energy contribution ($E_{C}$) and strong + kinetic energies contribution $E_{S}$:
$\omega=E_{C}+E_{S}=8.36\ \text{eV}\,.$ (23)
The sensitivity of the nuclear transition energy to changes in the fine structure constant $\alpha$ is
$\frac{\delta\omega}{\omega}=K_{\alpha}\frac{\delta\alpha}{\alpha}\,$, (24)
where $K_{\alpha}$ may be determined by the relation [50]
$K_{\alpha}=\frac{E_{C}}{\omega}\,.$ (25)
In Ref. [63] the sensitivity coefficient $K_{\alpha}$ was expressed via the difference in squared nuclear radii $\delta r^{2}$ and the difference of the electric quadrupole moments $\delta Q$ between the ground and isomeric states. Due to the lack of accurate data for electric quadrupole moments, Ref. [64] used the approximation of a constant nuclear density to express $\delta Q$ via a measured value of $\delta r^{2}$ and obtained $K_{\alpha}$ = $0.82\ (0.25)\times 10^{4}$ (with only the experimental error being presented). However, the approximation of a constant nuclear density is not sufficiently accurate in the case where the contributions of $\delta r^{2}$ and $\delta Q$ have opposing signs. Recent, more accurate measurements of $\delta r^{2}$ and $\delta Q$ [65], using the formula expressing $K_{\alpha}$ in terms of these parameters [63], give $K_{\alpha}$ = $0.59\ (0.23)\times 10^{4}$.
We may use this information, along with the results presented above, in order to estimate the nuclear transition’s sensitivity to changes in the QCD scale $\Lambda_{\text{QCD}}$. In similar fashion to the analysis performed in Refs. [50, 26], Equation (23) implies $E_{C}\approx-E_{S}$ and
$\frac{\delta\omega}{\omega}\approx\frac{E_{C}}{\omega}\left(\frac{\delta E_{C}}{E_{C}}-\frac{\delta E_{S}}{E_{S}}\right)\,.$ (26)
The Coulomb interaction contribution $E_{C}$ is proportional to $e^{2}/r_{0}\propto{\alpha}\Lambda_{\text{QCD}}$, see Eq. (16). Using Eq. (26), along with the sensitivity of $E_{S}$ to $\Lambda_{\text{QCD}}$ presented in (21), we yield
[…]
which implies that the nuclear clock transition is also more sensitive to variations in $\Lambda_{\text{QCD}}$.
This estimate for the sensitivity of the nuclear clock transition to variations of $\Lambda_{\text{QCD}}$ may have implications for the measurements performed in Ref. [66], which present constraints on the fractional variation of $\Lambda_{\text{QCD}}$ due to an interaction with ultralight dark matter, taking a rough estimate of the sensitivity coefficient as $K_{\text{QCD}}\sim K_{\alpha}\sim 10^{4}-10^{5}$. Eq. (27) allows one to refine such constraints, which are expected to be significantly improved in future measurements.
This effect is supposed to be extracted from the measurement of the variation of the dimensionless ratio of the nuclear transition frequency to an atomic transition frequency, which is proportional to the atomic unit of energy $E_{h}$. For example, in Ref. [60] the ratio of the nuclear transition frequency to an electron transition frequency in the Sr atom has been measured. The relative variation of the Sr transition frequency is not enhanced, so we may neglect its variation.
III Constraint from Big Bang Nucleosynthesis
Big-Bang Nucleosynthesis (BBN) took place during the first few minutes after the Big Bang, when the light nuclides H, D, He and Li were produced. Their primordial abundances are exquisitely sensitive to nuclear binding energies - and hence to the hadronic parameters that set those bindings. By comparing the observed primordial abundances with BBN calculations one can therefore bound any change in the underlying constants since the early Universe.
Ref. [43] (see previous estimates in Refs. [24, 25, 67, 68] and also [69]) computed how the binding energies in light nuclei respond to shifts of the hadron masses and, assuming the QCD scale $\Lambda_{\text{QCD}}$ is fixed, translated the difference between calculated and observed abundances into a fractional change of the average light-quark mass: $\delta m_{q}/m_{q}=K(0.013\pm 0.002)$, where $K\sim 1$ reflects a theoretical uncertainty of about a factor of 2.
Here we reinterpret that result in the complementary scenario in which the variation is driven instead by a gluon–scalar coupling that makes $\Lambda_{\text{QCD}}$ drift, while $m_{q}$ is held fixed. Using the nuclear binding energy response in Eq. (21), we obtain
$\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}=K(-2.5\pm 0.4)\times 10^{-3}\,.$ (28)
The fact that the central value is non-zero reflects the well-known “lithium problem” - the predicted 7Li abundance exceeds the observed abundance by a factor of 3 to 4. It seems that the lithium anomaly survives every Standard model update to date, and the variation of $\Lambda_{\text{QCD}}$ may solve this problem. The most credible fix within the Standard model is gradual Li destruction in old stars, supported by modern diffusion-mixing simulations but not yet universally accepted. Reviews of the primordial lithium problem - its observational status, nuclear physics inputs and proposed solutions - can be found in Refs. [70, 71].
IV Constraint from the Oklo natural nuclear reactor
The Oklo natural nuclear reactor was a naturally occurring fission reactor that operated about 1.8 billion years ago. Its sustained chain reactions produced a neutron spectrum that covered a very low energy resonance in 149Sm. This resonance, with $E_{r}=0.1\ \mathrm{eV}$, is small compared to the depth of the neutron potential (U=50MeVU=-50\ $\mathrm${MeV}), giving the system a high sensitivity to variations of the fundamental constants. Refs. [72, 73, 74, 75, 76] used this high sensitivity to find limits on the variation of $\alpha$, while Refs. [24, 25, 26] estimated limits on variation of the quark mass.
We can express the neutron capture resonance $E_{r}$ as
$\begin{split}E_{r}&=E_{k}+U,\\ E_{k}&=C_{1}\Lambda_{\text{QCD}}+C_{2}{\alpha}\Lambda_{\text{QCD}}+C_{3}m_{q},\\ U&=C_{4}\Lambda_{\text{QCD}}+C_{5}{\alpha}\Lambda_{\text{QCD}}+C_{6}m_{q},\end{split}$ (29)
where $E_{k}$ is the kinetic energy of the captured neutron inside the nucleus and $C_{i}$ are dimensionless parameters. The contribution of the Coulomb interaction is proportional to $e^{2}/r_{0}\propto{\alpha}\Lambda_{\text{QCD}}$, see Eq. (16). Substituting the expressions for $E_{k}$ and $U$ into the resonance energy gives
$E_{r}=(C_{1}+C_{4})\Lambda_{\text{QCD}}+(C_{2}+C_{5}){\alpha}\Lambda_{\text{QCD}}+(C_{3}+C_{6})m_{q}\,.$ (30)
Now, the condition $E_{k}=-U+0.1$eV implies that $E_{k}$ and $U$ are almost equal in magnitude. However, this near cancellation does not mean that the sensitivity of $E_{r}$ to the variation of fundamental constants is small, as the fact that $E_{r}\approx 0$ allows for all coefficients in Eq. (30) to be large, due to the fact that $E_{k}$ and $U$ are large.
Note that if only one parameter was involved (e.g. if we neglect the small contributions to the energy from $\alpha$ and $m_{q}$), the sensitivity to the variation of $\Lambda_{\text{QCD}}$ disappears, since $E_{k}$ and $U$ in this case are proportional ($E_{k}=CU$) and subsequently there will be no enhancement of the variation effect in their sum: $\delta E_{r}/E_{r}=\delta(E_{k}+U)/(E_{k}+U)=\delta E_{k}/E_{k}$. Another way one may consider this point is to note that the condition $E_{r}\approx 0$ would require $C_{1}+C_{4}\approx 0$ in the case that $m_{q}=\alpha=0$. As a result, the dependence on $\Lambda_{\text{QCD}}$ is eliminated in accordance with Eq. (30).
The experimental condition $E_{r}\approx 0$ allows us to express $C_{1}+C_{4}$ in terms of other constants. Doing so, we obtain
$\begin{split}E_{r}&=-\frac{\Lambda_{\text{QCD}}}{\Lambda^{0}_{\text{QCD}}}\bigg{[}(C_{2}+C_{5}){\alpha^{0}}\Lambda^{0}_{\text{QCD}}+(C_{3}+C_{6})m^{0}_{q}\bigg{]}\\ &+(C_{2}+C_{5}){\alpha}\Lambda_{\text{QCD}}+(C_{3}+C_{6})m_{q}\,.\end{split}$ (31)
where $\Lambda^{0}_{\text{QCD}}$, $\alpha^{0}$ and $m^{0}_{q}$ are the present values of these constants. In doing so, we obtain the following for the variation of the resonance energy
$\begin{split}\delta E_{r}&=-\frac{\delta\Lambda_{\text{QCD}}}{\Lambda^{0}_{\text{QCD}}}\bigg{[}(C_{2}+C_{5}){\alpha^{0}}\Lambda^{0}_{\text{QCD}}+(C_{3}+C_{6})m^{0}_{q}\bigg{]}\\ &+(C_{2}+C_{5})\delta(\alpha\Lambda_{\text{QCD}})+(C_{3}+C_{6})\delta m_{q}\,.\end{split}$ (32)
The shift of the resonance due to the variation of $\alpha$ was calculated in Refs. [72, 73] to be
$\delta E_{r}=(1.1\pm 0.1)\frac{\delta\alpha}{\alpha}\ \text{MeV}\,$, (33)
while the shift due to the variation of $m_{q}$
$\delta E_{r}\approx 10\frac{\delta m_{q}}{m_{q}}\ \text{MeV}\,$, (34)
was calculated in Ref. [26]. Using these results, we obtain
$\begin{split}\delta E_{r}&\approx 10\,{\text{MeV}}\,\left(-1.1\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}+\frac{\delta m_{q}}{m_{q}}-0.1\frac{\delta\alpha}{\alpha}\right)\,,\\ &\approx 10\,{\text{MeV}}\,\left(\frac{\delta X_{q}}{X_{q}}-0.1\frac{\delta\alpha}{\alpha}\right)\,.\end{split}$ (35)
Now we should specify how the shift of $E_{r}$ has been measured [72, 73]. It was expressed through the measured ratio of abundances of 149Sm and 150Sm near the Oklo natural nuclear reactor and compared to this ratio in other areas where there was no neutron flux transforming 149Sm to 150Sm. The neutron flux in Oklo was produced by the nuclear fission of 235U, and the integrated neutron flux was found by comparing the difference between the abundance of 235U in Oklo and its natural abundance in other places.
The observed shift in the resonance energy $\delta E_{r}$ between the present and time of the Oklo natural reactor 1.8 billion years ago was constrained to be less than 0.02 eV [75, 76, 74], which provides the constraint
$\bigg{|}\frac{\delta X_{q}}{X_{q}}-0.1\frac{\delta\alpha}{\alpha}\bigg{|}<2\times 10^{-9}\,.$ (36)
The variations of $\alpha$, $m_{q}$ and $\Lambda_{\text{QCD}}$ are produced by different terms in the interaction of dark matter or dark energy fields with gluons, quarks and photons. If we assume independent variations of these constants, we obtain limits on variation of these parameters during the 1.8 billion years
$\bigg{|}\frac{\delta\alpha}{\alpha}\bigg{|}$ $<2\times 10^{-8}\,$, (37)
$\bigg{|}\frac{\delta m_{q}}{m_{q}}\bigg{|}$ $<2\times 10^{-9}\,$, (38)
$\bigg{|}\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}\bigg{|}$ $<2\times 10^{-9}\,$, (39)
which may be compared to atomic clock results, by assuming the linear drift of these parameters:
$\bigg{|}\frac{\delta\alpha}{\alpha}\bigg{|}<1\times 10^{-17}\ \text{yr}^{-1}\,$, (40)
$\bigg{|}\frac{\delta m_{q}}{m_{q}}\bigg{|}<1\times 10^{-18}\ \text{yr}^{-1}\,$, (41)
$\bigg{|}\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}\bigg{|}<1\times 10^{-18}\ \text{yr}^{-1}\,.$ (42)
This chain of arguments is possible since the fission properties of 235U have a relatively low sensitivity to the variation of fundamental constants (compared to 149Sm).
This result can be interpreted in the context of dark-energy-type models [20, 21, 22, 23, 1], where variations occur on cosmological scales. Another class of models involves dark halos, in which the dark matter density around massive bodies is strongly enhanced compared to the average dark matter density and evolves on timescales of billions of years (see, e.g., [77, 78, 79]). Dynamical mechanisms of this type predict present-day fractional variations at the level of $10^{-13}$$10^{-16}$ [66], far below the $10^{-9}$ sensitivity accessible with Oklo. Therefore, Oklo may only probe an exceptionally large increase of the dark matter density. Large variations in the dark matter density could also accur due to passing dark matter clouds or Bose stars. In this context, Oklo constraints on $\Lambda_{\text{QCD}}$ may rule out the possibility that such extreme local enhancements of the dark matter density occurred in the past or are occurring today.
Now, we discuss the possible inconsistencies of this approach and seek a clarification. The results are based on the measurements of the dimensionless abundances, meaning they may be expressed in terms of dimensionless ratios of the fundamental constants, without the use of any “human-introduced” units like MeV (which also may be subject to variation) for the Coulomb and strong interaction’s contributions to energy, see Eq. (35). The key point is that we are interested in a resonance effect enhanced by a factor of $10^{8}$. As such, the choice of units is not important, as variations in the ratios of different units are not enhanced.
We still may pose a purely theoretical question, is it possible to avoid using “human-introduced” units and measure the effect in natural units for this problem from the start? The origin of the MeV coefficient in the expressions for the Coulomb and strong interaction energies can be traced back to $\Lambda_{\text{QCD}}$, which mainly determines nuclear properties. For example, the Coulomb interaction in nuclei is proportional to $e^{2}/r_{0}\sim\alpha\Lambda_{\text{QCD}}$. Therefore, we may make the substitution 10 MeV $\to\Lambda_{\text{QCD}}/30$. The change in the abundance of 149Sm is proportional to the change in the abundance of 235U, which produced the neutron flux captured by 149Sm. The proportionality coefficient includes the ratio of the resonance neutron capture cross section to a geometric factor, which is proportional to the squared internuclear distance in solids, approximately $a_{B}^{2}=1/(\alpha m_{e})^{2}$. Following this line of reasoning, we may conclude that the relevant dimensionless parameter is $E_{r}/(\alpha m_{e})$.
Further analysis, including addressing the neutron cooling process needed for an efficient neutron capture, involves a dependence on other fundamental physical constants such as Newton’s gravitational constant $G_{N}$, to which the Earth’s temperature is sensitive. However, these gravitational and other non-resonant effects are not enhanced and may be ignored.
V Constraints from Quasar Absorption Spectra
Quasar absorption spectra provide an extremely sensitive astrophysical probe of the variation of fundamental constants over cosmological timescales - see e.g [20, 21, 22, 23, 1, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96]. Intervening clouds along the line of sight of the quasar give access to the spectra of the atoms and ions present in the cloud. Calculations of the dependence of corresponding atomic transition frequencies on the fundamental constants have been performed e.g. in Refs. [97, 98, 14, 15, 99, 100, 101] and references therein. Through the analysis of atomic and molecular transitions in such absorption systems, measurements of the variation of the fine structure constant $\alpha$, the proton to electron mass ratio $\mu$ and composite parameters $F\equiv g_{p}(\alpha^{2}\mu)^{1.57}$, $x\equiv\alpha^{2}g_{p}/\mu$, $F^{\prime}\equiv\alpha^{2}\mu$ and $G\equiv g_{p}(\alpha\mu)^{1.85}$ have been performed, where $g_{p}$ is the proton $g$-factor; see e.g. Refs. [82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96].
Constant Constraint $z$ Ref. Lookback Time $\delta\Lambda_{\text{QCD}}/\Lambda_{\text{QCD}}$ $\delta\Lambda_{\text{QCD}}/\Lambda_{\text{QCD}}$
$(\times 10^{-5})$ $(\times 10^{10}\ \text{yr})$ ($\times 10^{-5}$) $(\times 10^{-15}\ \text{yr}^{-1})$
$\mu$ $(2.78\pm 0.88)$ 2.59 [82] 1.12 $(2.78\pm 0.88)$ $2.49\pm 0.79$
$(2.06\pm 0.79)$ 3.02 [82] 1.13 $(2.06\pm 0.79)$ $1.78\pm 0.68$
$(1.01\pm 0.62)$ 2.59 [83] 1.12 $(1.01\pm 0.62)$ $0.90\pm 0.55$
$(0.82\pm 0.74)$ 2.8 [83] 1.14 $(0.82\pm 0.74)$ $0.72\pm 0.65$
$(0.26\pm 0.30)$ 3.02 [83] 1.16 $(0.26\pm 0.30)$ $0.22\pm 0.26$
$(0.70\pm 0.80)$ 3.02 [84] 1.16 $(0.70\pm 0.80)$ $0.61\pm 0.69$
$(0.56\pm 0.55_{\text{stat}}\pm 0.27_{\text{syst}})$ 2.06 [85] 1.05 $(0.56\pm 0.61)$ $0.53\pm 0.58$
<0.18<0.18 0.685 [86] 0.629 <0.18<0.18 <0.29<0.29
<0.38<0.38 0.685 [87] 0.629 <0.38<0.38 <0.60<0.60
<0.14<0.14 0.89 [88] 0.638 <0.14<0.14 <0.22<0.22
<9<9 2.42 [89] 1.10 <9<9 <8.2<8.2
$F$ $(-0.44\pm 0.36\pm 1.0_{\text{syst}})$ 0.765 [90] 0.673 $(-0.28\pm 0.68)$ $0.42\pm 1.0$
$(0.51\pm 1.26)$ 0.247 [91] 0.293 $(0.33\pm 0.80)$ $1.11\pm 2.74$
$x$ $(-0.63\pm 0.99)$ 1.29 [92] 0.883 $(0.63\pm 0.99)$ $0.71\pm 1.12$
$(-0.17\pm 0.17)$ 3.17 [93] 1.17 $(0.17\pm 0.17)$ $0.15\pm 0.15$
$F^{\prime}$ $(1.0\pm 10)$ 4.69 [94] 1.24 $(1.00\pm 10.00)$ $0.80\pm 8.0$
$(14\pm 15)$ 6.42 [94] 1.29 $(14\pm 15)$ $11\pm 12$
$G$ <1.1<1.1 0.247 [95] 0.294 <0.60<0.60 <2.03<2.03
<1.16<1.16 0.0018 [95] 0.0025 <0.63<0.63 <250<250
$(-1.18\pm 0.46)$ 0.247 [96] 0.294 $(-0.64\pm 0.25)$ $2.17\pm 0.85$
Table 3: Constraints on the variation of $\Lambda_{\text{QCD}}$ from Quasar Absorption Spectra. The constants (as presented in Ref. [1]) are: $\mu\equiv m_{p}/m_{e}$, $F\equiv g_{p}(\alpha^{2}\mu)^{1.57}$, $x\equiv\alpha^{2}g_{p}/\mu$, $F^{\prime}\equiv\alpha^{2}\mu$ and $G\equiv g_{p}(\alpha\mu)^{1.85}$.
Aside from $\alpha$, these parameters are all sensitive to changes in $\Lambda_{\text{QCD}}$. In this section, we re-interpret constraints on the variation of these $\Lambda_{\text{QCD}}$-dependent parameters from such measurements as limits on the variation of $\Lambda_{\text{QCD}}$. We once again do so under the assumption that this parameter acquires a time dependence upon interaction with an ultralight dark matter or dark energy field which couples to the gluon field, leaving $\alpha$ and the electron mass unchanged.
The proton mass $m_{p}$ and magnetic $g$-factor $g_{p}$ both depend on $\Lambda_{\text{QCD}}$. Within 10% accuracy, $\frac{\delta\mu}{\mu}\approx\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}$ - see Eq. (20). A detailed calculation of the quark mass dependence of $g_{p}$ in Ref. [14] together with Eqs. (5,6) leads to the estimate $\frac{\delta g_{p}}{g_{p}}\approx 0.10\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}$. This allows us to translate existing astrophysical bounds, as compiled in Ref. [1], into direct constraints on $\Lambda_{\text{QCD}}$, see Table 3. Under the assumption of a linear variation of $\Lambda_{\text{QCD}}$, these constraints are remarkably close to those obtained from atomic clock measurements. We note, however, that this assumption has no justification over such an extensive period of time.
VI Summary
Some theories beyond the Standard model predict that “fundamental constants” might drift (or oscillate) with time or space. Only dimensionless constants - such as the fine-structure constant $\alpha$ can be constrained in a model-independent way: any change in a pure number is unambiguous. By contrast, limits on dimensionful quantities ($c$, $\hbar$, $G_{N}$, $\Lambda_{\text{QCD}}$, $m_{e}$,…) make sense only within a specified framework - for example, a model where an ultralight dark-matter or dark-energy field couples to ordinary matter. The reason is simple: the “human” units used to quote those quantities (seconds, metres, electron-volts, Hertz, …) are themselves defined through physical processes that depend on the underlying constants. Thus, the units may vary too and their dependence on the fundamental constants may be quite complicated. For example, the dependence of the hyperfine transition frequency in Cs, which defines the second and the Hertz, can be seen in Table 1. Meaningful constraints on dimensionful parameters therefore require stating how the units are anchored in the chosen model; otherwise one can always absorb an apparent variation by redefining the ruler or the clock.
To consider the variation of dimensionful constants one should specify an interaction which produces this variation. In this paper, we have investigated the sensitivity of several systems to the variation of the dimensionful QCD interaction constant $\Lambda_{\text{QCD}}$ and extracted limits on its drift. In particular, we consider the case in which an interaction between an evolving dark matter or dark energy scalar field $\phi$ and the gluon field induces a time dependence in $\Lambda_{\text{QCD}}$, while leaving $\alpha$ and the fundamental masses unchanged.
Within the minimal scenario in which this gluonic channel is the sole $\phi$ interaction, we show that the atomic Yb/Cs clock comparison [4] places the constraint $\dot{\Lambda}_{\text{QCD}}/\Lambda_{\text{QCD}}=(3.2\pm 3.5)\times 10^{-17}\ \text{yr}^{-1}$. Further independent constraints are obtained from isotopic ratios measured in the Oklo natural nuclear reactor, which limit the variation to $|\delta\Lambda_{\text{QCD}}/\Lambda_{\text{QCD}}|<2\times 10^{-9}$ over 1.8 billion years. Assuming a linear drift, this translates to a bound $|\dot{\Lambda}_{\text{QCD}}/\Lambda_{\text{QCD}}|<1\times 10^{-18}\text{yr}^{-1}$. Additionally, the sensitivity of the 229Th nuclear clock to the variation of $\Lambda_{\text{QCD}}$ is estimated to be four orders of magnitude higher than that of current atomic clocks.
We have also obtained constraints on $\Lambda_{\text{QCD}}$ variation from quasar absorption spectra (see Table 3) and Big Bang Nucleosynthesis data. Comparison of these astrophysical and cosmological constraints with atomic clock limits requires the assumption of a linear variation of $\Lambda_{\text{QCD}}$, which has no justification over such large time intervals (from 3 to 13 billion years). However, if a linear variation is assumed, limits derived from the quasar absorption spectra are surprisingly close to those from atomic clocks.
The BBN data indicates a non-zero deviation from the present value, $\delta\Lambda_{\text{QCD}}/\Lambda_{\text{QCD}}=(-2.5\pm 0.4)\times 10^{-3}$. This is a reflection of the well-known “lithium problem,” wherein the predicted abundance of 7Li within the Standard Model exceeds the observed abundance by a factor of 3 to 4. The variation of $\Lambda_{\text{QCD}}$ may fix this problem. However, the search for solutions within the Standard Model continues.
Acknowledgements
We are grateful to Dima Budker, Abhishek Banerjee and Gilad Perez for valuable discussions and for bringing relevant references to our attention. This work was supported by the Australian Research Council Grant No. DP230101058.

### v4

@@
 We also obtain constraints from quasar absorption spectra and Big Bang Nucleosynthesis data.
-Introduction
+I Introduction
 The search for the variation of fundamental constants is a well-established area of research motivated by theories beyond the Standard model that unify gravity with the other fundamental forces, as well as cosmological models involving scalar fields which predict a space-time variation of these constants.
 Constraints on such variation have been derived from Big Bang Nucleosynthesis (BBN), quasar absorption spectra, the cosmic microwave background (CMB) and the Oklo natural nuclear reactor, see e.g.
-the review [1] and references therein.
+the review [Uzan] and references therein.
 The stringent bounds on the linear drift of the fundamental constants have been obtained using measurements of the temporal drift of atomic clock transition frequencies.
@@
 As a result, these measurements may be used to place constraints on the variation of these constants.
-Experimental results have been obtained by measuring the oscillating frequency ratios of electron transitions in a range of systems, including Dy/Cs [2], Rb/Cs [3], Yb/Cs [4], Sr/H/Si cavity [5], Cs/cavity [6], Cs/H [7], Al+/Hg+ [8] and Yb+/Yb+/Sr [9, 10].
+Experimental results have been obtained by measuring the oscillating frequency ratios of electron transitions in a range of systems, including Dy/Cs [DyCs], Rb/Cs [Hees2016], Yb/Cs [YbCs], Sr/H/Si cavity [HSi], Cs/cavity [Tretiak], Cs/H [Fischer2004], Al+/Hg+ [AlHgdrift] and Yb+/Yb+/Sr [Banerjee2023, Filzinger2023].
 The results of these experiments have been used to place constraints on the variations of fundamental constants using the sensitivity calculations performed in Refs.
-[11, 12, 13, 14, 15, 16, 17] and several other papers referenced there.
+[PRLWebb, PRAWebb, CanJPh, Thomas, Tedesco, csquarks, Borschevsky] and several other papers referenced there.
 It is important to note that if the absolute measurement of a transition frequency is performed, the dimensionless ratio of this frequency to the Cs hyperfine transition frequency, which defines the units of frequency and time, is being measured.
-The Cs transition frequency has a complicated dependence on the fundamental constants [15].
+The Cs transition frequency has a complicated dependence on the fundamental constants [Tedesco].
 This principle applies to any measurement of variations of fundamental constants in any field of physics and astrophysics.
@@
 Consequently, interpreting experimental results in terms of changing dimensionful constants is not model-independent; one must instead identify a set of dimensionless fundamental constants relevant to this measurement and calculate the dependence of an observed effect on these constants  [^1] It was noted in Refs.
-[18, 19], that the number of possible ”dimensionless” constants may increase when we allow for fast variations of the constants, where ”fast” is determined by the time scale of the response of the studied species or experimental apparatus used.
+[Antypas1, Antypas2], that the number of possible ”dimensionless” constants may increase when we allow for fast variations of the constants, where ”fast” is determined by the time scale of the response of the studied species or experimental apparatus used.
 In this case, the relevant dimensionless quantity is, for example, the ratio $m_{e}/\expectationvalue{m_{e}}$, and $\expectationvalue{m_{e}}$ is the time average.
@@
 To be consistent, one must include every relevant term in the Lagrangian when deriving how the dimensionless ratios of fundamental constants vary for a given measurement.
-Alternatively, one may speak about the apparent variation of the fundamental constants, which may be just a convenient way to calculate the effects of specific terms in the interaction of Standard model particles with a dark matter (or dark energy [20, 21, 22, 23, 1]) scalar or pseudoscalar field $\phi$.
+Alternatively, one may speak about the apparent variation of the fundamental constants, which may be just a convenient way to calculate the effects of specific terms in the interaction of Standard model particles with a dark matter (or dark energy [BarrowMagueijo1998, Barrow2005, BarrowSandvikMagueijo2002, Wetterich2003, Uzan]) scalar or pseudoscalar field $\phi$.
 For example, the interaction of fermions with $\phi$, which may be presented in the form $g\phi^{n}m\bar{f}f$, is similar to the fermion mass term $m\bar{f}f$.
@@
 The effects of the variation of fermion masses and the variation of $\alpha$ have been calculated; see, e.g.
-[11, 12, 13, 24, 25, 14, 15, 16, 17].
+[PRLWebb, PRAWebb, CanJPh, Shuryak2002, FlambaumShuryak, Thomas, Tedesco, csquarks, Borschevsky].
 Note that in this approach, the variation of measurement units is still a relevant issue if one considers measurements of dimensionful parameters.
@@
 It is usually assumed that the comparison of two optical clock transitions is sensitive to variations of the fine structure constant only; however, following Ref.
-[9], we argue that this system is also sensitive to variations of $\Lambda_{\text{QCD}}$ and repurpose existing results to place constraints.
+[Banerjee2023], we argue that this system is also sensitive to variations of $\Lambda_{\text{QCD}}$ and repurpose existing results to place constraints.
 We also examine the sensitivity of the fractional variation of the transition frequency between the ground state and the low-energy isomer in 229Th to variations in the QCD scale $\Lambda_{\text{QCD}}$, which appears to be enhanced relative to the proposed nuclear clock’s sensitivity to the quark mass $m_{q}$.
@@
 Finally, we reinterpret existing constraints from the Oklo natural nuclear reactor to place constraints on the variation of $\Lambda_{\text{QCD}}$.
-I Constraints on $\Lambda_{\text{QCD}}$ via the Reinterpretation of Clock constraints
+II Constraints on $\Lambda_{\text{QCD}}$ from atomic clocks measurements
+II.1 Ratios of magnetic dipole hyperfine and optical transition frequencies
 The experimental observation of variation in nature may be interpreted in a model-independent manner as the variation of at least one of three independent fundamental dimensionless ratios;
@@
 where $X_{q,e}$ corresponds to the ratio of the quark/electron mass to the QCD scale $\Lambda_{\text{QCD}}$.
-Here, we are working in natural units, setting $c=1$ and assuming that $\Lambda_{\text{QCD}}$ has dimension of mass.
+Here, we are working in natural units, $c=\hbar=1$.
 For example, the variation of the ratio of two atomic, molecular or nuclear transition frequencies may be presented as
@@
 In the previous calculations in Refs.
-[14, 15], the dependence on $X_{q}$ was calculated assuming that $\Lambda_{\text{QCD}}$ is constant and only the quark masses vary.
+[Thomas, Tedesco], the dependence on $X_{q}$ was calculated assuming that $\Lambda_{\text{QCD}}$ is constant and only the quark masses vary.
 Below, we will demonstrate that there is no need for such an assumption.
@@
 The dependence of the nuclear magnetic moments on $m_{q}$ was calculated in Refs.
-[14, 15].
+[Thomas].
 A less significant contribution to the $m_{q}$ dependence, which originates from the variation of the nuclear radius, was calculated in Refs.
-[26, 27].
+[Wiringa2009, Dinh2009].
 The nuclear magnetic moment and nuclear radius are also sensitive to the QCD scale $\Lambda_{\text{QCD}}$.
@@
 These corrections provide a natural mechanism by which these magnetic moments can acquire a dependence on both the quark mass and $\Lambda_{\text{QCD}}$.
-The nucleon magnetic moment including such corrections can be schematically written as [28, 29]:
+The nucleon magnetic moment including such corrections can be schematically written as [Bernard1995, Kubis2001]:
 $\mu_{N}=\mu_{N}^{(0)}+\mu_{N,\pi}^{(1)}\,$, (4)
-where $\mu_{N}^{(0)}$ denotes the static low-energy constant contributions at tree-level and $\mu_{N,\pi}^{(1)}$ represents the dominant corrections to the nucleon magnetic moments, measured in units of the nuclear magneton $e\hbar/2m_{p}c$, which at leading one-loop order are proportional to the pion mass $m_{\pi}$ [30, 31]:
+where $\mu_{N}^{(0)}$ denotes the static low-energy constant contributions at tree-level and $\mu_{N,\pi}^{(1)}$ represents the dominant corrections to the nucleon magnetic moments, measured in units of the nuclear magneton $e\hbar/2m_{p}c$, which at leading one-loop order are proportional to the pion mass $m_{\pi}$ [Bernard1992, Jenkins1993]:
 $\mu_{N,\pi}^{(1)}\sim\frac{g_{A}^{2}m_{\pi}m_{N}}{8\pi f_{\pi}^{2}}\,$, (5)
@@
 The sign of this correction is positive for the proton magnetic moment and negative for the neutron magnetic moment.
-Since $m_{\pi}^{2}\sim m_{q}\Lambda_{\text{QCD}}$ via the Gell-Mann–Oakes–Renner relation  [32] and $f_{\pi}\propto\Lambda_{\text{QCD}}$, the pion mass inherits a dependence on the light quark mass $m_{q}$ and $\Lambda_{\text{QCD}}$.
+Since $m_{\pi}^{2}\sim m_{q}\Lambda_{\text{QCD}}$ via the Gell-Mann–Oakes–Renner relation [GellMan1968] and $f_{\pi}\propto\Lambda_{\text{QCD}}$, the pion mass inherits a dependence on the light quark mass $m_{q}$ and $\Lambda_{\text{QCD}}$.
 Consequently, the one-loop correction scales as:
@@
 Clock 1 Clock 2 Constraint (yr-1) Ref. Constants dependence
-87Rb 133 Cs $(-0.5\pm 5.3)\times 10^{-16}$ [33] $\alpha^{-0.49}X_{q}^{-0.025}$
-1H 133Cs $(-32\pm 63)\times 10^{-16}$ [34] $\alpha^{2.83}X_{q}^{-0.039}X_{e}$
-199Hg+ 133Cs $(3.7\pm 3.9)\times 10^{-16}$ [35] $\alpha^{6.05}X_{q}^{-0.039}X_{e}$
-171Yb+ 133Cs $(-3.1\pm 3.4)\times 10^{-17}$ [36] $\alpha^{1.93}X_{q}^{-0.039}X_{e}$
-87Sr 133Cs $(-1.0\pm 1.8)\times 10^{-15}$ [37] $\alpha^{2.77}X_{q}^{-0.039}X_{e}$
+87Rb 133 Cs $(-1.4\pm 0.9)\times 10^{-16}$ [RbCsNew] $\alpha^{-0.49}X_{q}^{-0.025}$
+1H 133Cs $(-32\pm 63)\times 10^{-16}$ [FischerHCs2004] $\alpha^{2.83}X_{q}^{-0.039}X_{e}$
+199Hg+ 133Cs $(3.7\pm 3.9)\times 10^{-16}$ [Fortier2007] $\alpha^{6.05}X_{q}^{-0.039}X_{e}$
+171Yb+ 133Cs $(-3.1\pm 3.4)\times 10^{-17}$ [Lange2021] $\alpha^{1.93}X_{q}^{-0.039}X_{e}$
+87Sr 133Cs $(-1.0\pm 1.8)\times 10^{-15}$ [Blatt2008] $\alpha^{2.77}X_{q}^{-0.039}X_{e}$
 Table 1: Constraints on the variation of the ratio of clock frequencies.
@@
 Under the minimal assumption that this gluonic channel is the sole $\phi$ interaction, the constraints presented in Table 1 can be translated into limits on the temporal variation of the QCD scale, $\Lambda_{\text{QCD}}$, via the dependence of each transition on $X_{q}$ and $X_{e}$.
-The dominant contribution to the sensitivity arises from the $X_{e}$ term, and the tightest constraints may be obtained from the 171Yb+/Cs clock measurement [36]:
+The dominant contribution to the sensitivity arises from the $X_{e}$ term, and the tightest constraints may be obtained from the 171Yb+/Cs clock measurement [Lange2021]:
 $\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}=(3.2\pm 3.5)\times 10^{-17}\ \text{yr}^{-1}\,.$ (7)
-I.1 Sensitivity of the comparison of two optical clock transition frequencies to $X_{e}$
+II.2 Ratio of electric quadrupole hyperfine to magnetic dipole hyperfine and optical transition frequencies
+In our paper [FlambaumMansour2023] we found that the electric quadrupole hyperfine constant $B$ has higher sensitivity to variation of $m_{q}$ than the magnetic hyperfine constant $A$, and suggested to measure variation of the ratio $B/A$.
+In this subsection we will show that this ratio $B/A$, as well as the ratio of $B$ to an optical transition frequency, are highly sensitive to the variation of $\Lambda_{\text{QCD}}$.
+The dependence of the hyperfine constants $B$ and $A$ on the fundamental constants is given by:
+$B$ $\propto\frac{eQ}{a_{B}^{3}}R_{B}(Z\alpha)\sim\frac{e^{2}r_{0}^{2}}{a_{B}^{3}}R_{B}(Z\alpha)\,$, (8)
+$A$ $\propto\frac{\mu_{B}\mu_{N}}{a_{B}^{3}}R_{A}(Z\alpha)\sim\frac{e^{2}\hbar^{2}g_{N}}{m_{e}m_{N}c^{2}a_{B}^{3}}R_{A}(Z\alpha)\,$, (9)
+where $r_{0}\approx 1.2\ \text{fm}$ is the internucleon distance, $a_{B}$ is the Bohr radius, $\mu_{B}$ and $\mu_{N}$ are the Bohr and nuclear magnetons and $g_{N}$ is the nuclear magnetic $g$-factor, $Q$ is the electric quadrupole moment, $R_{B}(Z\alpha)$ and $R_{A}(Z\alpha)$ are the relativistic factors for the electric quadrupole and magnetic dipole constants respectively, which are presented e.g.
+in the paper [Tedesco] and the book [Sobelman].
+Therefore, the ratio of the electric quadrupole and magnetic dipole constants may be written as
+$\frac{B}{A}\propto\frac{r_{0}^{2}m_{e}m_{N}c^{2}R_{B}(Z\alpha)}{g_{N}\hbar^{2}R_{A}(Z\alpha)}\,.$ (10)
+The dominating effect in the variation of this ratio comes from the variation of $r_{0}^{2}$ and $m_{N}$.
+The dependence of $g_{N}$ on hadronic parameters is generally weak and varies from nucleus to nucleus [Thomas, Tedesco].
+To leading order, this dependence can be neglected, allowing for general estimates that are valid across all nuclei.
+The variation of the internucleon distance $r_{0}$ in terms of the pion and nucleon mass has been calculated in Ref.
+[Wiringa2009]  [^2] The dependence on the nucleon and pion masses arises mainly from the kinetic energy and the long-distance behavior of the strong interaction.
+These contributions can be calculated reliably and are included in the present work.
+Ref.
+[Wiringa2009] also considered vector meson and $\Delta$ contributions, which describe short-distance nucleon interaction.
+The description of the short- distance interaction is model-dependent.
+Also, the $V$ and $\Delta$ contributions partially cancel.
+Since their inclusion does not clearly improve the accuracy, we omit the $V$ and $\Delta$ contributions.:
+$\frac{\delta r_{0}}{r_{0}}=1.8\frac{\delta m_{\pi}}{m_{\pi}}-4.8\frac{\delta m_{N}}{m_{N}}\,.$ (11)
+The variation of the nucleon mass $m_{N}$ can be related to the variation of the fundamental constants by [Thomas, Damour2010, SHIFMAN1978, Junnarkar2013, Aoki2020, Oswald2022]
+[…]
+where $m_{q,-}=(m_{u}-m_{d})$ is the mass difference.
+The dependence of the variation of the nucleon mass on this parameter is weak, and we may neglect it.
+We will also assume that $\frac{\delta m_{s}}{m_{s}}=\frac{\delta m_{q}}{m_{q}}$, where $m_{s}$ is the mass of the strange quark (this assumption may be motivated by the Higgs mechanism of mass generation).
+Noting the dependence of the pion mass on $\Lambda_{\text{QCD}}$ and $m_{q}$ from the Gell-Mann-Oakes-Renner relation [GellMan1968]
+$m_{\pi}^{2}\sim\Lambda_{\text{QCD}}m_{q}\,$, (13)
+we yield the following dependence on the hadronic parameters
+$\frac{\delta r_{0}}{r_{0}}=-3.5\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}+0.3\frac{\delta m_{q}}{m_{q}}\,.$ (14)
+Thus, using Eqs. (10) and (14), we yield
+$\frac{\delta(B/A)}{(B/A)}=-6.0\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}+0.7\frac{\delta m_{q}}{m_{q}}\,.$ (15)
+It may also be useful to estimate the dependence of the ratio of the electric quadrupole constant $B$ to the optical transition energy $E_{\text{opt}}\propto e^{2}/a_{B}$. Using Eq. (8) we obtain
+$\frac{\delta(B/E_{\text{opt}})}{(B/E_{opt})}\approx\frac{\delta(r_{0}^{2}/a_{B}^{2})}{r_{0}^{2}/a_{B}^{2}}.$ (16)
+We omit here the relativistic factors which contribute to the $\alpha$-dependence. The dependence on hadronic parameters may be presented as
+$\frac{\delta(B/E_{\text{opt}})}{(B/E_{opt})}=-9.6\frac{\delta m_{N}}{m_{N}}+3.6\frac{\delta m_{\pi}}{m_{\pi}}$ (17)
+$=-6.9\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}+0.6\frac{\delta m_{q}}{m_{q}}.$ (18)
+II.3 Sensitivity of the comparison of two optical clock transition frequencies to hadronic parameters
 Different types of atomic transition frequencies have a differing dependence on the fundamental constants.
@@
 Following Ref.
-[9], we aim to propose a similar dependence on $X_{e}$ for the ratio of two optical frequencies, by probing the contributions of nuclear mass $m_{N}$ and nuclear radius $r_{N}$ to atomic transition frequencies (these contributions produce mass and field isotope shifts).
+[Banerjee2023], we aim to propose a similar dependence on $X_{e}$ for the ratio of two optical frequencies, by probing the contributions of nuclear mass $m_{N}$ and nuclear radius $r_{N}$ to atomic transition frequencies (these contributions produce mass and field isotope shifts).
 These contributions are small when compared to the optical transition frequency; however, this mechanism indicates a sensitivity to the parameter $X_{e}=m_{e}/\Lambda_{\text{QCD}}$ from measurements comparing the ratio of two optical transition frequencies.
@@
 $K_{4}$ $=\frac{K_{\text{MS}}^{\nu_{b}}}{\nu_{b}m_{N,b}}\,$, (23)
-where $K_{\text{MS}}$ and $K_{\text{FS}}$ are the mass and field shift coefficients respectively and $m_{N}$ is the mass of an atom with atomic mass number $A$ (which is largely determined by the nuclear mass $m_{N}$).
+where $K_{\text{MS}}$ and $K_{\text{FS}}$ are the conventional mass and field shift coefficients respectively and $m_{N}$ is the mass of the nucleus.
 The mass shift term dominates for light nuclei whilst the field shift term dominates for heavy nuclei.
-Results have been obtained for the ratio of the optical transition frequencies of $^{1}S_{0}\rightarrow\ ^{3}P_{0}$ transitions in Al+ and $^{2}S_{1/2}\rightarrow\ ^{2}D_{5/2}$ transitions in Hg+ [38, 39, 40] and the ratio of the $^{2}S_{1/2}(F=0)\leftrightarrow\ ^{2}F_{7/2}(F=3)$ electric octupole (E3) and $^{2}S_{1/2}(F=0)\leftrightarrow\ ^{2}D_{3/2}(F=2)$ electric quadrupole (E2) transition frequencies in Yb+ [9].
-We have previously used Equation (8) along with these results to place constraints on the variation of hadronic parameters, see Refs.
-[41, 40].
-When considering low-mass nuclei, the sensitivity to the parameter $X_{e}$ may be yielded from the dependence of optical transition frequencies on the reduced mass (the normal mass shift)
-$\bar{\mu}=\frac{m_{e}m_{N}}{m_{e}+m_{N}}\,$, (13)
-where the nuclear mass $m_{N}$ is proportional to the nucleon mass, $m_{N}\approx Am_{n}$. The sensitivity to other hadronic parameters comes about from the energy shift due to a finite nuclear size.
+Before proceeding to specific numbers, we discuss what dimensionless parameters are involved in these variations.
 These effects are extracted from the measurement of the variation of the ratio of two atomic transition frequencies which have different dependencies on varying parameters.
 In this case, one may say that the dimensionless parameters appear after dividing the variations of dimensionful parameters by the appropriate atomic units of energy $E_{h}=\alpha^{2}m_{e}c^{2}$, length $a_{B}=\hbar/(mc\alpha)$, etc.
+The nucleon mass dependence in light nuclei originates from the dependence of optical transition frequencies on the reduced mass  [^3] This equation accounts for the normal mass shift only, the specific mass shift depends on the transition and is omitted here.
+$\bar{\mu}=\frac{m_{e}m_{N}}{m_{e}+m_{N}}\,$, (24)
+where the nuclear mass $m_{N}$ is proportional to the nucleon mass, $m_{N}\approx Am_{n}$. Therefore, the dimensionless parameter here is $X_{e}=m_{e}/m_{n}\approx m_{e}/\Lambda_{\text{QCD}}$.
+The sensitivity to other hadronic parameters comes about from the energy shift due to a finite nuclear size.
 In high-mass nuclei, the relative change in energy (or electronic transition frequency $\nu$) in atoms is proportional to
 $\frac{\delta E}{E}\propto\left(\frac{r_{N}}{a_{B}}\right)^{2}=\left(\frac{r_{N}m_{e}\alpha c}{\hbar}\right)^{2}\,.$ (25)
-Setting $c=\hbar=1$, and noting the fact that the nuclear radii in all nuclei are reasonably described by a universal formula $r_{N}=A^{1/3}r_{0}$, where $r_{0}\approx 1.2\ \text{fm}$ is the internucleon distance, we yield
+Setting $r_{N}=A^{1/3}r_{0}$, we yield
 $\left(\frac{r_{N}}{a_{B}}\right)^{2}=\left(r_{0}A^{1/3}m_{e}\alpha\right)^{2}\,.$ (26)
-The distance between the nucleons $r_{0}$ is determined by the minimum of the potential in the nucleon-nucleon interaction.
-In the Walecka model [42] this potential is produced by a vector meson exchange repulsion at short distances and sigma meson exchange attraction at larger distances.
-Both the vector meson and sigma meson masses are proportional to $\Lambda_{\text{QCD}}$, so the range of these potentials is inversely proportional to $\Lambda_{\text{QCD}}$.
-Thus $r_{0}\propto\frac{1}{\Lambda_{\text{QCD}}}$ and
-$\left(\frac{r_{N}}{a_{B}}\right)^{2}$ $\propto\left(\frac{A^{1/3}m_{e}\alpha}{\Lambda_{\text{QCD}}}\right)^{2}$ (16)
-Assuming independent variations, the constraints on the variation of the nuclear charge radius $r_{N}$ presented in these references may be repurposed into constraints on the parameter $X_{e}$, noting that the variation of the $r_{N}^{2}$ is proportional to (see Equation (16))
-$\frac{\delta(r_{N}/a_{B})^{2}}{(r_{N}/a_{B})^{2}}\sim 2\left(\frac{\delta m_{e}}{m_{e}}+\frac{\delta\alpha}{\alpha}-\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}\right)\sim 2\left(\frac{\delta X_{e}}{X_{e}}+\frac{\delta\alpha}{\alpha}\right)\,.$ (17)
-The numerical constraints obtained from the systems of interest are presented in Table 2.
+The variation of $r_{N}^{2}$ is proportional to (see Eqns (26) and (14))
+$\frac{\delta(r_{N}/a_{B})^{2}}{(r_{N}/a_{B})^{2}}$ $=2\left(\frac{\delta m_{e}}{m_{e}}+\frac{\delta\alpha}{\alpha}-3.5\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}+0.3\frac{\delta m_{q}}{m_{q}}\right)\,.$ (27)
+Now we discuss specific numerical results which have been derived from the ratio of the optical transition frequencies of $^{1}S_{0}\rightarrow\ ^{3}P_{0}$ transitions in Al+ and $^{2}S_{1/2}\rightarrow\ ^{2}D_{5/2}$ transitions in Hg+ [Wineland2007, AlHgdrift, Tang2021, DzubaFlambaumMansour2024] and the ratio of the $^{2}S_{1/2}(F=0)\leftrightarrow\ ^{2}F_{7/2}(F=3)$ electric octupole (E3) and $^{2}S_{1/2}(F=0)\leftrightarrow\ ^{2}D_{3/2}(F=2)$ electric quadrupole (E2) transition frequencies in Yb+ [Filzinger2023, Banerjee2023].
+In our papers Refs.
+[FlambaumMansour2023, DzubaFlambaumMansour2024] we used these results to obtain constrains on variation of nucleon mass, nuclear radius, and other hadron parameters.
+The constraints presented in these references and Eq.
+(27) may be repurposed into constraints on $\Lambda_{\text{QCD}}$.
+These constraints are presented in Table 2.
 Clock 1 Clock 2 Constraint (yr-1) Ref. $\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}$ (yr-1)
-Al+ Hg+ $(-5.3\pm 7.9)\times 10^{-17}$ [8] $(5.5\pm 8.5)\times 10^{-15}$
-Yb+(E3) Yb+(E2) $(-1.2\pm 1.8)\times 10^{-18}$ [10] $(0.25\pm 0.38)\times 10^{-15}$
+Al+ Hg+ $(-5.3\pm 7.9)\times 10^{-17}$ [AlHgdrift] $(3.2\pm 4.9)\times 10^{-15}$
+Yb+(E3) Yb+(E2) $(-1.2\pm 1.8)\times 10^{-18}$ [Filzinger2023] $(7.2\pm 11)\times 10^{-17}$
 Table 2: Constraints on fractional variation of $\Lambda_{\text{QCD}}$ from the comparison of two optical clock transition frequencies in different systems.
-II Sensitivity of the 229Th based nuclear clock to $\Lambda_{\text{QCD}}$
+III Sensitivity of the 229Th based nuclear clock to $\Lambda_{\text{QCD}}$
 Refs.
-[43, 26] have calculated the sensitivity of the binding energy and nuclear radius in various light nuclei to the different hadron masses.
+[Wiringa2007, Wiringa2009] have calculated the sensitivity of the binding energy and nuclear radius in various light nuclei to the different hadron masses.
 These calculations were extrapolated to heavier nuclei.
@@
 Using Table III of Ref.
-[26], the variation of the binding energy $E_{S}$ is related to the variation of the hadron masses by
+[Wiringa2009], the variation of the binding energy $E_{S}$ is related to the variation of the hadron masses by  [^4] The contribution of the electromagnetic interaction will be considered below.
 $\frac{\delta E_{S}}{E_{S}}\approx\Delta\mathcal{E}(m_{N})\frac{\delta m_{N}}{m_{N}}+\Delta\mathcal{E}(m_{\pi})\frac{\delta m_{\pi}}{m_{\pi}}\,$, (28)
@@
 which were roughly calculated to be $\Delta\mathcal{E}(m_{N})\approx 12$ and $\Delta\mathcal{E}(m_{\pi})\approx-6$.
-The variation of the nucleon mass $m_{N}$ can be related to the variation of the fundamental constants by [14, 44, 45, 46, 47, 48]
-[…]
-where $m_{q,-}=(m_{u}-m_{d})$ is the mass difference.
-The dependence of the variation of the nucleon mass on this parameter is weak, and we may neglect it.
-We will also assume that $\frac{\delta m_{s}}{m_{s}}=\frac{\delta m_{q}}{m_{q}}$, where $m_{s}$ is the mass of the strange quark (this assumption may be motivated by the Higgs mechanism of mass generation).
-Noting the dependence of the pion mass on $\Lambda_{\text{QCD}}$ and $m_{q}$ from the Gell-Mann-Oakes-Renner relation, $m_{\pi}^{2}\sim\Lambda_{\text{QCD}}m_{q}$ [32], we yield the following result for the variation of the binding energy $E_{S}$
+Using Eqs.
+(12) and (13), we yield the following result for the variation of the binding energy $E_{S}$
 $\frac{\delta E_{S}}{E_{S}}\approx 7.9\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}-1.5\frac{\delta m_{q}}{m_{q}}\,.$ (30)
@@
 However, this expression does not include the contribution of the electromagnetic energy $E_{C}$ which is small in light nuclei ($E_{C}\propto Z(Z-1)$).
-It may be instructive to estimate separately kinetic energy contribution $E_{K}\propto 1/m_{n}r_{N}^{2}$.
-Using Equation (20), together with $r_{N}=A^{1/3}r_{0}$ and the scaling $r_{0}\propto\Lambda_{\text{QCD}}^{-1}$, the fractional variation is
-[…]
-implying that in the kinetic energy the $\Lambda_{\text{QCD}}$ term strongly dominates over the $m_{q}$ term.
 These results have implications for experiments with the “nuclear clock” transition between the ground and first excited states of the 229Th nucleus, first proposed in Ref.
-[49], where effects of the variation of fundamental constants are strongly enhanced [50].
+[Peik2003], where effects of the variation of fundamental constants are strongly enhanced [FlambaumTh2006].
 This transition has been measured by a range of experiments, see e.g.
-[51, 52, 53, 54, 55, 56, 57, 58, 59, 60].
-The most precise measurement gives a transition frequency of $\omega=2,020,407,384,335\ (2)$ kHz ( 8.36 eV) for 229Th dopant ions embedded in a calcium fluoride crystal [60].
+[Seiferle2019, Masuda2019, Yamaguchi2019, Sikorski2020, Katori2024, ZhangBao2024, Tiedau2024, Elwell2024, Hiraki2024, Zhang2024].
+The most precise measurement gives a transition frequency of $\omega=2,020,407,384,335\ (2)$ kHz ( 8.36 eV) for 229Th dopant ions embedded in a calcium fluoride crystal [Zhang2024].
 Using the results presented above, we may write this transition frequency $\omega$ as a sum of nearly canceling components that have different dependencies on the fundamental constants.
-The low-lying isomeric state in 229Th is theorised to exhibit a different shape and spin configuration to the ground state [61, 62], resulting in a small energy shift due to the residual strong and electromagnetic effects.
+The low-lying isomeric state in 229Th is theorised to exhibit a different shape and spin configuration to the ground state [Litvinova, Minkov2021], resulting in a small energy shift due to the residual strong and electromagnetic effects.
 As such, we may write $\omega$ as a sum of the electromagnetic energy contribution ($E_{C}$) and strong + kinetic energies contribution $E_{S}$:
@@
 $\frac{\delta\omega}{\omega}=K_{\alpha}\frac{\delta\alpha}{\alpha}\,$, (32)
-where $K_{\alpha}$ may be determined by the relation [50]
+where $K_{\alpha}$ may be determined by the relation [FlambaumTh2006]
 $K_{\alpha}=\frac{E_{C}}{\omega}\,.$ (33)
 In Ref.
-[63] the sensitivity coefficient $K_{\alpha}$ was expressed via the difference in squared nuclear radii $\delta r^{2}$ and the difference of the electric quadrupole moments $\delta Q$ between the ground and isomeric states.
-Due to the lack of accurate data for electric quadrupole moments, Ref.
-[64] used the approximation of a constant nuclear density to express $\delta Q$ via a measured value of $\delta r^{2}$ and obtained $K_{\alpha}$ = $0.82\ (0.25)\times 10^{4}$ (with only the experimental error being presented).
+[PhysRevLett.102.210801] the sensitivity coefficient $K_{\alpha}$ was expressed via the difference in squared nuclear radii $\delta r^{2}$ and the difference of the electric quadrupole moments $\delta Q$ between the ground and isomeric states.
+[^5] Due to the lack of accurate data for electric quadrupole moments, Ref.
+[Fadeev2020] used the approximation of a constant nuclear density to express $\delta Q$ via a measured value of $\delta r^{2}$ and obtained $K_{\alpha}$ = $0.82\ (0.25)\times 10^{4}$, with only the experimental error being presented.
 However, the approximation of a constant nuclear density is not sufficiently accurate in the case where the contributions of $\delta r^{2}$ and $\delta Q$ have opposing signs.
-Recent, more accurate measurements of $\delta r^{2}$ and $\delta Q$ [65], using the formula expressing $K_{\alpha}$ in terms of these parameters [63], give $K_{\alpha}$ = $0.59\ (0.23)\times 10^{4}$.
+Recent, measurements of $\delta r^{2}$ and $\delta Q$ [SafronovaYe], using the formula expressing $K_{\alpha}$ in terms of these parameters [PhysRevLett.102.210801], give $K_{\alpha}$ = $0.59\ (0.23)\times 10^{4}$  [^6] Ref.
+[SafronovaYe] also investigated the sensitivity of $K_{\alpha}$ to a possible octupole deformation.
+.
 We may use this information, along with the results presented above, in order to estimate the nuclear transition’s sensitivity to changes in the QCD scale $\Lambda_{\text{QCD}}$.
 In similar fashion to the analysis performed in Refs.
-[50, 26], Equation (23) implies $E_{C}\approx-E_{S}$ and
+[FlambaumTh2006, Wiringa2009], Equation (31) implies $E_{C}\approx-E_{S}$ and
 $\frac{\delta\omega}{\omega}\approx\frac{E_{C}}{\omega}\left(\frac{\delta E_{C}}{E_{C}}-\frac{\delta E_{S}}{E_{S}}\right)\,.$ (34)
-The Coulomb interaction contribution $E_{C}$ is proportional to $e^{2}/r_{0}\propto{\alpha}\Lambda_{\text{QCD}}$, see Eq.
-(16).
-Using Eq.
-(26), along with the sensitivity of $E_{S}$ to $\Lambda_{\text{QCD}}$ presented in (21), we yield
+The Coulomb interaction contribution $E_{C}$ is proportional to $e^{2}/r_{0}\propto{\alpha}\Lambda_{\text{QCD}}$, see Eq. (27).
+Using Eq. (34), along with the sensitivity of $E_{S}$ to $\Lambda_{\text{QCD}}$ presented in (30), we yield
 […]
@@
 This estimate for the sensitivity of the nuclear clock transition to variations of $\Lambda_{\text{QCD}}$ may have implications for the measurements performed in Ref.
-[66], which present constraints on the fractional variation of $\Lambda_{\text{QCD}}$ due to an interaction with ultralight dark matter, taking a rough estimate of the sensitivity coefficient as $K_{\text{QCD}}\sim K_{\alpha}\sim 10^{4}-10^{5}$.
+[Fuchs2024PRX], which present constraints on the fractional variation of $\Lambda_{\text{QCD}}$ due to an interaction with ultralight dark matter, taking a rough estimate of the sensitivity coefficient as $K_{\text{QCD}}\sim K_{\alpha}\sim 10^{4}-10^{5}$.
 Eq.
@@
 For example, in Ref.
-[60] the ratio of the nuclear transition frequency to an electron transition frequency in the Sr atom has been measured.
+[Zhang2024] the ratio of the nuclear transition frequency to an electron transition frequency in the Sr atom has been measured.
 The relative variation of the Sr transition frequency is not enhanced, so we may neglect its variation.
-III Constraint from Big Bang Nucleosynthesis
+IV Constraint from Big Bang Nucleosynthesis
 Big-Bang Nucleosynthesis (BBN) took place during the first few minutes after the Big Bang, when the light nuclides H, D, He and Li were produced.
@@
 Ref.
-[43] (see previous estimates in Refs.
-[24, 25, 67, 68] and also [69]) computed how the binding energies in light nuclei respond to shifts of the hadron masses and, assuming the QCD scale $\Lambda_{\text{QCD}}$ is fixed, translated the difference between calculated and observed abundances into a fractional change of the average light-quark mass: $\delta m_{q}/m_{q}=K(0.013\pm 0.002)$, where $K\sim 1$ reflects a theoretical uncertainty of about a factor of 2.
+[Wiringa2007] (see previous estimates in Refs.
+[Shuryak2002, FlambaumShuryak, Dmitriev2003, Dmitriev2004] and also [BERENGUT2010114]) computed how the binding energies in light nuclei respond to shifts of the hadron masses and, assuming the QCD scale $\Lambda_{\text{QCD}}$ is fixed, translated the difference between calculated and observed abundances into a fractional change of the average light-quark mass: $\delta m_{q}/m_{q}=K(0.013\pm 0.002)$, where $K\sim 1$ reflects a theoretical uncertainty of about a factor of 2.
 Here we reinterpret that result in the complementary scenario in which the variation is driven instead by a gluon–scalar coupling that makes $\Lambda_{\text{QCD}}$ drift, while $m_{q}$ is held fixed.
@@
 Reviews of the primordial lithium problem - its observational status, nuclear physics inputs and proposed solutions - can be found in Refs.
-[70, 71].
-IV Constraint from the Oklo natural nuclear reactor
+[Fields2011, Cyburt2016].
+V Constraint from the Oklo natural nuclear reactor
 The Oklo natural nuclear reactor was a naturally occurring fission reactor that operated about 1.8 billion years ago.
@@
 Refs.
-[72, 73, 74, 75, 76] used this high sensitivity to find limits on the variation of $\alpha$, while Refs.
-[24, 25, 26] estimated limits on variation of the quark mass.
+[Shlyakhter1976, DAMOUR1996, FUJII2000, Gould2006, Petrov2006] used this high sensitivity to find limits on the variation of $\alpha$, while Refs.
+[Shuryak2002, FlambaumShuryak, Wiringa2009] estimated limits on variation of the quark mass.
 We can express the neutron capture resonance $E_{r}$ as
@@
 where $E_{k}$ is the kinetic energy of the captured neutron inside the nucleus and $C_{i}$ are dimensionless parameters.
-The contribution of the Coulomb interaction is proportional to $e^{2}/r_{0}\propto{\alpha}\Lambda_{\text{QCD}}$, see Eq.
-(16).
+The contribution of the Coulomb interaction is proportional to $e^{2}/r_{0}\propto{\alpha}\Lambda_{\text{QCD}}$.
 Substituting the expressions for $E_{k}$ and $U$ into the resonance energy gives
@@
 $\begin{split}\delta E_{r}&=-\frac{\delta\Lambda_{\text{QCD}}}{\Lambda^{0}_{\text{QCD}}}\bigg[(C_{2}+C_{5}){\alpha^{0}}\Lambda^{0}_{\text{QCD}}+(C_{3}+C_{6})m^{0}_{q}\bigg]\\ &+(C_{2}+C_{5})\delta(\alpha\Lambda_{\text{QCD}})+(C_{3}+C_{6})\delta m_{q}\,.\end{split}$ (40)
-The shift of the resonance due to the variation of $\alpha$ was calculated in Refs. [72, 73] to be
+The shift of the resonance due to the variation of $\alpha$ was calculated in Refs. [Shlyakhter1976, DAMOUR1996] to be
 $\delta E_{r}=(1.1\pm 0.1)\frac{\delta\alpha}{\alpha}\ \text{MeV}\,$, (41)
@@
 $\delta E_{r}\approx 10\frac{\delta m_{q}}{m_{q}}\ \text{MeV}\,$, (42)
-was calculated in Ref. [26]. Using these results, we obtain
+was calculated in Ref. [Wiringa2009]. Using these results, we obtain
 $\begin{split}\delta E_{r}&\approx 10\,{\text{MeV}}\,\left(-1.1\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}+\frac{\delta m_{q}}{m_{q}}-0.1\frac{\delta\alpha}{\alpha}\right)\,,\\ &\approx 10\,{\text{MeV}}\,\left(\frac{\delta X_{q}}{X_{q}}-0.1\frac{\delta\alpha}{\alpha}\right)\,.\end{split}$ (43)
-Now we should specify how the shift of $E_{r}$ has been measured [72, 73].
+Now we should specify how the shift of $E_{r}$ has been measured [Shlyakhter1976, DAMOUR1996].
 It was expressed through the measured ratio of abundances of 149Sm and 150Sm near the Oklo natural nuclear reactor and compared to this ratio in other areas where there was no neutron flux transforming 149Sm to 150Sm.
 The neutron flux in Oklo was produced by the nuclear fission of 235U, and the integrated neutron flux was found by comparing the difference between the abundance of 235U in Oklo and its natural abundance in other places.
-The observed shift in the resonance energy $\delta E_{r}$ between the present and time of the Oklo natural reactor 1.8 billion years ago was constrained to be less than 0.02 eV [75, 76, 74], which provides the constraint
+The observed shift in the resonance energy $\delta E_{r}$ between the present and time of the Oklo natural reactor 1.8 billion years ago was constrained to be less than 0.02 eV [Gould2006, Petrov2006, FUJII2000], which provides the constraint
 $\bigg|\frac{\delta X_{q}}{X_{q}}-0.1\frac{\delta\alpha}{\alpha}\bigg|<2\times 10^{-9}\,.$ (44)
@@
 This chain of arguments is possible since the fission properties of 235U have a relatively low sensitivity to the variation of fundamental constants (compared to 149Sm).
-This result can be interpreted in the context of dark-energy-type models [20, 21, 22, 23, 1], where variations occur on cosmological scales.
-Another class of models involves dark halos, in which the dark matter density around massive bodies is strongly enhanced compared to the average dark matter density and evolves on timescales of billions of years (see, e.g., [77, 78, 79]).
-Dynamical mechanisms of this type predict present-day fractional variations at the level of $10^{-13}$$10^{-16}$ [66], far below the $10^{-9}$ sensitivity accessible with Oklo.
+This result can be interpreted in the context of dark-energy-type models [BarrowMagueijo1998, Barrow2005, BarrowSandvikMagueijo2002, Wetterich2003, Uzan], where variations occur on cosmological scales.
+Another class of models involves dark halos, in which the dark matter density around massive bodies is strongly enhanced compared to the average dark matter density and evolves on timescales of billions of years (see, e.g., [BanerjeeFlambaum2020, Banerjee2020, Budker2023]).
+Dynamical mechanisms of this type predict present-day fractional variations at the level of $10^{-13}$$10^{-16}$ [Fuchs2024PRX], far below the $10^{-9}$ sensitivity accessible with Oklo.
 Therefore, Oklo may only probe an exceptionally large increase of the dark matter density.
@@
 However, these gravitational and other non-resonant effects are not enhanced and may be ignored.
-V Constraints from Quasar Absorption Spectra
-Quasar absorption spectra provide an extremely sensitive astrophysical probe of the variation of fundamental constants over cosmological timescales - see e.g [20, 21, 22, 23, 1, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96].
+VI Constraints from Quasar Absorption Spectra
+Quasar absorption spectra provide an extremely sensitive astrophysical probe of the variation of fundamental constants over cosmological timescales - see e.g [BarrowMagueijo1998, Barrow2005, BarrowSandvikMagueijo2002, Wetterich2003, Uzan, Murphy2001, MurphyWebbFlambaum2003, Reinhold2006, King2008, Thompson2009, Malec2009, Murphy2008, Menten2008, Henkel2009, Petitjean2009, Kanekar2005, Darling2004, Tzanavaris2007, Srianand2010, Levshakov2008, Chengalur2003, Kanekar2010].
 Intervening clouds along the line of sight of the quasar give access to the spectra of the atoms and ions present in the cloud.
@@
 in Refs.
-[97, 98, 14, 15, 99, 100, 101] and references therein.
+[DZUBAPRL1999, DZUBAPRA1999, Thomas, Tedesco, CanJPhQ, Ammonia, KozlovMolecules] and references therein.
 Through the analysis of atomic and molecular transitions in such absorption systems, measurements of the variation of the fine structure constant $\alpha$, the proton to electron mass ratio $\mu$ and composite parameters $F\equiv g_{p}(\alpha^{2}\mu)^{1.57}$, $x\equiv\alpha^{2}g_{p}/\mu$, $F^{\prime}\equiv\alpha^{2}\mu$ and $G\equiv g_{p}(\alpha\mu)^{1.85}$ have been performed, where $g_{p}$ is the proton $g$-factor; see e.g.
 Refs.
-[82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96].
-Constant Constraint $z$ Ref. Lookback Time $\delta\Lambda_{\text{QCD}}/\Lambda_{\text{QCD}}$ $\delta\Lambda_{\text{QCD}}/\Lambda_{\text{QCD}}$
-$(\times 10^{-5})$ $(\times 10^{10}\ \text{yr})$ ($\times 10^{-5}$) $(\times 10^{-15}\ \text{yr}^{-1})$
-$\mu$ $(2.78\pm 0.88)$ 2.59 [82] 1.12 $(2.78\pm 0.88)$ $2.49\pm 0.79$
-$(2.06\pm 0.79)$ 3.02 [82] 1.13 $(2.06\pm 0.79)$ $1.78\pm 0.68$
-$(1.01\pm 0.62)$ 2.59 [83] 1.12 $(1.01\pm 0.62)$ $0.90\pm 0.55$
-$(0.82\pm 0.74)$ 2.8 [83] 1.14 $(0.82\pm 0.74)$ $0.72\pm 0.65$
-$(0.26\pm 0.30)$ 3.02 [83] 1.16 $(0.26\pm 0.30)$ $0.22\pm 0.26$
-$(0.70\pm 0.80)$ 3.02 [84] 1.16 $(0.70\pm 0.80)$ $0.61\pm 0.69$
-$(0.56\pm 0.55_{\text{stat}}\pm 0.27_{\text{syst}})$ 2.06 [85] 1.05 $(0.56\pm 0.61)$ $0.53\pm 0.58$
-<0.18<0.18 0.685 [86] 0.629 <0.18<0.18 <0.29<0.29
-<0.38<0.38 0.685 [87] 0.629 <0.38<0.38 <0.60<0.60
-<0.14<0.14 0.89 [88] 0.638 <0.14<0.14 <0.22<0.22
-<9<9 2.42 [89] 1.10 <9<9 <8.2<8.2
-$F$ $(-0.44\pm 0.36\pm 1.0_{\text{syst}})$ 0.765 [90] 0.673 $(-0.28\pm 0.68)$ $0.42\pm 1.0$
-$(0.51\pm 1.26)$ 0.247 [91] 0.293 $(0.33\pm 0.80)$ $1.11\pm 2.74$
-$x$ $(-0.63\pm 0.99)$ 1.29 [92] 0.883 $(0.63\pm 0.99)$ $0.71\pm 1.12$
-$(-0.17\pm 0.17)$ 3.17 [93] 1.17 $(0.17\pm 0.17)$ $0.15\pm 0.15$
-$F^{\prime}$ $(1.0\pm 10)$ 4.69 [94] 1.24 $(1.00\pm 10.00)$ $0.80\pm 8.0$
-$(14\pm 15)$ 6.42 [94] 1.29 $(14\pm 15)$ $11\pm 12$
-$G$ <1.1<1.1 0.247 [95] 0.294 <0.60<0.60 <2.03<2.03
-<1.16<1.16 0.0018 [95] 0.0025 <0.63<0.63 <250<250
-$(-1.18\pm 0.46)$ 0.247 [96] 0.294 $(-0.64\pm 0.25)$ $2.17\pm 0.85$
-Table 3: Constraints on the variation of $\Lambda_{\text{QCD}}$ from Quasar Absorption Spectra.
-The constants (as presented in Ref.
-[1]) are: $\mu\equiv m_{p}/m_{e}$, $F\equiv g_{p}(\alpha^{2}\mu)^{1.57}$, $x\equiv\alpha^{2}g_{p}/\mu$, $F^{\prime}\equiv\alpha^{2}\mu$ and $G\equiv g_{p}(\alpha\mu)^{1.85}$.
+[Reinhold2006, King2008, Thompson2009, Malec2009, Murphy2008, Menten2008, Henkel2009, Petitjean2009, Kanekar2005, Darling2004, Tzanavaris2007, Srianand2010, Levshakov2008, Chengalur2003, Kanekar2010].
 Aside from $\alpha$, these parameters are all sensitive to changes in $\Lambda_{\text{QCD}}$.
@@
 A detailed calculation of the quark mass dependence of $g_{p}$ in Ref.
-[14] together with Eqs.
+[Thomas] together with Eqs.
 (5,6) leads to the estimate $\frac{\delta g_{p}}{g_{p}}\approx 0.10\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}$.
 This allows us to translate existing astrophysical bounds, as compiled in Ref.
-[1], into direct constraints on $\Lambda_{\text{QCD}}$, see Table 3.
+[Uzan], into direct constraints on $\Lambda_{\text{QCD}}$, see Table 3.
 Under the assumption of a linear variation of $\Lambda_{\text{QCD}}$, these constraints are remarkably close to those obtained from atomic clock measurements.
 We note, however, that this assumption has no justification over such an extensive period of time.
-VI Summary
+VII Summary
 Some theories beyond the Standard model predict that “fundamental constants” might drift (or oscillate) with time or space.
@@
 Meaningful constraints on dimensionful parameters therefore require stating how the units are anchored in the chosen model; otherwise one can always absorb an apparent variation by redefining the ruler or the clock.
+Constant Constraint $z$ Ref. Lookback Time $\delta\Lambda_{\text{QCD}}/\Lambda_{\text{QCD}}$ $\delta\Lambda_{\text{QCD}}/\Lambda_{\text{QCD}}$
+$(\times 10^{-5})$ $(\times 10^{10}\ \text{yr})$ ($\times 10^{-5}$) $(\times 10^{-15}\ \text{yr}^{-1})$
+$\mu$ $(2.78\pm 0.88)$ 2.59 [Reinhold2006] 1.12 $(2.78\pm 0.88)$ $2.49\pm 0.79$
+$(2.06\pm 0.79)$ 3.02 [Reinhold2006] 1.13 $(2.06\pm 0.79)$ $1.78\pm 0.68$
+$(1.01\pm 0.62)$ 2.59 [King2008] 1.12 $(1.01\pm 0.62)$ $0.90\pm 0.55$
+$(0.82\pm 0.74)$ 2.8 [King2008] 1.14 $(0.82\pm 0.74)$ $0.72\pm 0.65$
+$(0.26\pm 0.30)$ 3.02 [King2008] 1.16 $(0.26\pm 0.30)$ $0.22\pm 0.26$
+$(0.70\pm 0.80)$ 3.02 [Thompson2009] 1.16 $(0.70\pm 0.80)$ $0.61\pm 0.69$
+$(0.56\pm 0.55_{\text{stat}}\pm 0.27_{\text{syst}})$ 2.06 [Malec2009] 1.05 $(0.56\pm 0.61)$ $0.53\pm 0.58$
+<0.18<0.18 0.685 [Murphy2008] 0.629 <0.18<0.18 <0.29<0.29
+<0.38<0.38 0.685 [Menten2008] 0.629 <0.38<0.38 <0.60<0.60
+<0.14<0.14 0.89 [Henkel2009] 0.638 <0.14<0.14 <0.22<0.22
+<9<9 2.42 [Petitjean2009] 1.10 <9<9 <8.2<8.2
+$F$ $(-0.44\pm 0.36\pm 1.0_{\text{syst}})$ 0.765 [Kanekar2005] 0.673 $(-0.28\pm 0.68)$ $0.42\pm 1.0$
+$(0.51\pm 1.26)$ 0.247 [Darling2004] 0.293 $(0.33\pm 0.80)$ $1.11\pm 2.74$
+$x$ $(-0.63\pm 0.99)$ 1.29 [Tzanavaris2007] 0.883 $(0.63\pm 0.99)$ $0.71\pm 1.12$
+$(-0.17\pm 0.17)$ 3.17 [Srianand2010] 1.17 $(0.17\pm 0.17)$ $0.15\pm 0.15$
+$F^{\prime}$ $(1.0\pm 10)$ 4.69 [Levshakov2008] 1.24 $(1.00\pm 10.00)$ $0.80\pm 8.0$
+$(14\pm 15)$ 6.42 [Levshakov2008] 1.29 $(14\pm 15)$ $11\pm 12$
+$G$ <1.1<1.1 0.247 [Chengalur2003] 0.294 <0.60<0.60 <2.03<2.03
+<1.16<1.16 0.0018 [Chengalur2003] 0.0025 <0.63<0.63 <250<250
+$(-1.18\pm 0.46)$ 0.247 [Kanekar2010] 0.294 $(-0.64\pm 0.25)$ $2.17\pm 0.85$
+Table 3: Constraints on the variation of $\Lambda_{\text{QCD}}$ from Quasar Absorption Spectra.
+The constants (as presented in Ref.
+[Uzan]) are: $\mu\equiv m_{p}/m_{e}$, $F\equiv g_{p}(\alpha^{2}\mu)^{1.57}$, $x\equiv\alpha^{2}g_{p}/\mu$, $F^{\prime}\equiv\alpha^{2}\mu$ and $G\equiv g_{p}(\alpha\mu)^{1.85}$.
 To consider the variation of dimensionful constants one should specify an interaction which produces this variation.
@@
 In particular, we consider the case in which an interaction between an evolving dark matter or dark energy scalar field $\phi$ and the gluon field induces a time dependence in $\Lambda_{\text{QCD}}$, while leaving $\alpha$ and the fundamental masses unchanged.
-Within the minimal scenario in which this gluonic channel is the sole $\phi$ interaction, we show that the atomic Yb/Cs clock comparison [4] places the constraint $\dot{\Lambda}_{\text{QCD}}/\Lambda_{\text{QCD}}=(3.2\pm 3.5)\times 10^{-17}\ \text{yr}^{-1}$.
+Within the minimal scenario in which this gluonic channel is the sole $\phi$ interaction, we show that the atomic Yb/Cs clock comparison [YbCs] places the constraint $\dot{\Lambda}_{\text{QCD}}/\Lambda_{\text{QCD}}=(3.2\pm 3.5)\times 10^{-17}\ \text{yr}^{-1}$.
 Further independent constraints are obtained from isotopic ratios measured in the Oklo natural nuclear reactor, which limit the variation to $|\delta\Lambda_{\text{QCD}}/\Lambda_{\text{QCD}}|<2\times 10^{-9}$ over 1.8 billion years.